# Recursive Language Models

### 31 Dec 2025 v1

@@ 2. ACT: Take an action (execute code)
D.2 Summary agent baseline
@@
Appendix E Additional Benchmark Details
@@
E.1 OOLONG-Pairs Benchmark
To create OOLONG-Pairs, we synthetically generate $20$ new tasks based on the ground-truth labels for the OOLONG Bertsch et al. (2025) trec_coarse split for input contexts of length in [1024, 2048, 4096, 8192, 16384, 32768, 65536, 131072, 262144, 524288, 1048576]. Similar to OOLONG, each question requires correctly predicing the semantic mapping for each entry.
Ensuring $\mathcal{I}\approx O(N^{2})$ on OOLONG-Pairs. We noticed that many tasks that aggregate over pairs of entries could actually be solved without looking at the pairs and only looking at each entry in a linear fashion (e.g. using the principle of inclusion-exclusion in set theory), so we explicitly created questions that ask for all pairs satisfying some properties.
@@ E.1 OOLONG-Pairs Benchmark
E.2 Scaling Huge Document Corpuses in BrowseComp+

### 28 Jan 2026 v2

@@ 1. Use FINAL(your final answer here) to provide the answer directly
(1c) The diff of the system prompt for RLM with REPL (Qwen3-8B), which has a few changes from the GPT-5 prompt due to differences in context length and similar sub-calling behavior as Qwen3-Coder-480B-A35B:
--- a/REPL_SYSTEM_PROMPT.txt
+++ b/REPL_SYSTEM_PROMPT_QWEN3_8B.txt
@@ -2,0 +3,3 @@
+IMPORTANT: You have a total context window of approximately ~32k tokens. Be very careful about context length limits. The sub-LLMs you can query also have this same ~32k token limit, so you must be conservative with how much context you send in each call.
@@ -7 +10 @@
-2. A llm_query function that allows you to query an LLM (that can handle around 500K chars) inside your REPL environment.
+2. A llm_query function that allows you to query an LLM (that can handle around ~100k chars, roughly 32k tokens) inside your REPL environment.
@@ -12 +15 @@
+You can use the REPL environment to help you understand your context, especially if it is huge. Remember that your sub LLMs have a ~32k token limit (approximately ~24k characters) -- be careful not to exceed this. For example, a viable strategy is to feed 2-3 documents per sub-LLM query. Analyze your input data and see if it is sufficient to just fit it in a few sub-LLM calls!
+IMPORTANT: Be very careful about using llm_query as it incurs high runtime costs. Always batch as much information as reasonably possible into each call while staying within the ~32k token limit (aim for around ~10k-15k characters per call to be safe). For example, if you have 1000 lines of information to process, its much better to split into chunks of 50-100 and call llm_query on each chunk (10-20 calls total) rather than making 1000 individual calls. Minimize the number of llm_query calls by batching related information together, but always respect the ~32k token limit.
@@ -15 +20 @@
-chunk = context[:10000]
+chunk = context[:1000]
@@ -62,0 +68 @@
+FINAL_VAR(final_answer)
@@ -66 +73 @@
+IMPORTANT: When you are done with the iterative process, you MUST provide a final answer inside a FINAL function when you have completed your task, NOT in code or repl tags. Do not use these tags unless you have completed your task. You have two options:
@@ 2. ACT: Take an action (execute code)
C.2 Summary agent baseline
@@
Appendix D Additional Benchmark Details
@@
D.1 OOLONG-Pairs Benchmark
To create OOLONG-Pairs, we synthetically generate $20$ new tasks based on the ground-truth labels for the OOLONG (Bertsch et al., 2025) trec_coarse split for input contexts of length in [1024, 2048, 4096, 8192, 16384, 32768, 65536, 131072, 262144, 524288, 1048576]. Similar to OOLONG, each question requires correctly predicing the semantic mapping for each entry.
Ensuring quadratic scaling on OOLONG-Pairs. We noticed that many tasks that aggregate over pairs of entries could actually be solved without looking at the pairs and only looking at each entry in a linear fashion (e.g. using the principle of inclusion-exclusion in set theory), so we explicitly created questions that ask for all pairs satisfying some properties.
@@ D.1 OOLONG-Pairs Benchmark
D.2 Scaling Huge Document Corpuses in BrowseComp+
@@
Appendix E Additional RLM Trajectories
In this section, we provide several example trajectories to highlight characteristics of frontier models as RLMs. Many of the trajectories are too long to fit in text, so we describe each step and show specific examples when relevant.
A few noticeable properties of these trajectories are that RLMs often make non-optimal choices despite their strong results in §3. For example, in Example E.2, we observed that the RLM with Qwen3-Coder carefully constructs its final answer through a mix of recursive sub-calls and code execution in the first iteration, but then discards this information and continues wasting sub-calls before not using these stored answers. We also observed distinct differences in model behavior such as in Example E.3, where we found Qwen3-Coder make hundreds to thousands of recursive sub-calls for a single simple task, while GPT-5 makes on the order of ten. While these examples are not comprehensive, they provide useful qualitative insight into how to improve RLMs.
E.1 RLM(GPT-5) on BrowseComp-Plus-Query_74
E.2 RLM(Qwen3-Coder) on OOLONG-Pairs-Query_3
E.3 RLM(Qwen3-Coder) on OOLONG-Query_212
E.4 RLM(GPT-5) on CodeQA-Query_44
Final. The RLM answers choice ‘1’, which is the correct answer.
Appendix F Additional Runtime and Cost Analysis of RLMs
We supplement the cost and runtime analysis of RLMs with additional, fine-grained plots. In Figures 9, 10 we include a histogram for the cost of each method on every task for both GPT-5 and Qwen3-Coder. We generally observe long-tailed, high-variance trajectories for RLMs in both models.
Figure 7: Plotted quartiles of the runtime GPT-5 across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §3.2. We plot the 25th, 50th, 75th, and 95th percentiles.
Figure 8: Plotted quartiles of the runtime Qwen3-Coder-480B across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §3.2. We plot the 25th, 50th, 75th, and 95th percentiles.
Figure 9: Histogram of the API costs for GPT-5 across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §3.2.
Figure 10: Histogram of the API costs for Qwen3-Coder-480B across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §3.2.
Figure 11: We plot the API cost in USD for the runs in Figure 1.